# Efficient Graph Optimization via Distance-Aware Graph Representation Learning

### v6

@@ Abstract
We propose DRTR, a efficient graph optimization framework that integrates distance-aware multi-hop message passing with dynamic topology refinement. Unlike standard GNNs that rely on shallow, fixed-hop aggregation, DRTR leverages both static preprocessing and dynamic resampling to capture deeper structural dependencies. A Distance Recomputator prunes semantically weak edges using adaptive attention, while a Topology Reconstructor establishes latent connections among distant but relevant nodes. This joint mechanism enables more expressive and robust representation learning across evolving graph structures. Extensive experiments demonstrate that DRTR outperforms baseline GNNs in both accuracy and scalability, especially in complex and noisy graph environments.
@@ 1 Introduction
Graph Neural Networks (GNNs) have emerged as a powerful paradigm for semi-supervised node classification and graph representation learning, particularly due to their ability to propagate information over graph structures via message passing Kipf and Welling (2017); Hamilton et al. (2017). Classic GNNs operate by aggregating features from multi-hop neighborhoods using static architectures and fixed aggregation depths. However, such designs may fall short when handling graphs with noisy connections, uneven structural densities, or dynamically evolving topologies Liu et al. (2021); Wang et al. (2021).
Recent studies have started exploring adaptive multi-hop aggregation. For example, GAMLP Zhang et al. (2022) applies distinct MLPs for each hop-wise representation, while ImprovingTE Yao et al. (2023) enriches hop-specific encodings with contextual substructures. Nevertheless, these methods still depend heavily on fixed-hop sampling strategies, which may hinder generalization and limit the ability to capture complex or noisy graph signals.
In this work, we propose a novel adaptive reconstruction framework for multi-hop GNNs that dynamically adjusts both distance computations and local graph structures to promote more effective and robust message passing. Inspired by coreset selection Guo et al. (2022), our method adaptively reconstructs the neighborhood for each node, enabling the model to better capture latent semantic patterns and mitigate the propagation of noise.
@@
Unlike previous models that statically aggregate neighborhood features, DRTR continuously adapts its perception of graph topology and node relationships throughout training. This dynamic reconstruction mechanism not only improves the quality of learned representations but also enables better generalization across a wide range of graphs.
To validate the effectiveness of our framework, we conduct comprehensive experiments across several benchmark datasets, including both static and dynamic graph settings. Our results demonstrate that DRTR consistently outperforms state-of-the-art multi-hop GNN models in terms of accuracy, robustness, and stability. Furthermore, our theoretical analysis sheds light on why adaptive graph refinement enhances representation learning by aligning information flow with the underlying graph semantics.
In summary, our contributions are as follows:
  • 1.
    We propose DRTR, a novel adaptive reconstruction framework that dynamically refines node distances and topological structures to enhance multi-hop message passing.
  • 2.
    We design two complementary modules—the Distance Recomputator and the Topology Reconstructor—to improve representation quality and model robustness in both static and evolving graph scenarios.
  • 3.
    We provide both empirical and theoretical evidence showing that DRTR achieves superior performance over strong baselines while improving stability and adaptability in graph representation tasks.
@@ 2 Related Works
Conventional Graph Neural Networks (GNNs) such as GCN Kipf and Welling (2017), GraphSAGE Hamilton et al. (2017), and GAT Veličković et al. (2018) adopt fixed-hop message passing schemes and pre-defined neighbor aggregation strategies. However, recent studies recognize that structural connections in graphs often suffer from noise, imbalance, or redundancy Liu et al. (2023). To address this, various works have explored structure learning Franceschi et al. (2020); Chen et al. (2020); Liu et al. (2021), which attempts to refine or reconstruct the graph topology during training. Nonetheless, many such methods rely on end-to-end optimization or learn a static edge mask, which lacks the flexibility to respond to evolving node semantics during multi-hop propagation.
@@
Several studies have attempted to incorporate distance or importance-based filtering into message passing. For instance, Personalized PageRank (PPR)-based methods Gasteiger et al. (2019); Chien et al. (2021) reweight neighbor contributions based on precomputed transition probabilities. Meanwhile, GAMLP Zhang et al. (2022) and ImprovingTE Yao et al. (2023) explicitly leverage multi-hop propagation, with hop-specific encoders and contextual subgraph tokens. However, these approaches still depend on fixed-hop sampling, which may not fully capture varying node proximities or information decay across hops. Unlike prior work, DRTR dynamically recomputes node distances during training using the proposed Distance Recomputator, allowing for more context-aware and adaptive neighborhood construction.
@@
Traditional GNNs typically perform synchronous aggregation, where all neighbor messages are processed simultaneously within each hop Wu et al. (2021). This design overlooks the fact that different neighbors may carry drastically different relevance to the target node. Asynchronous strategies Li et al. (2017); Schaefer et al. (2022) attempt to stagger updates or introduce temporal dynamics, but often target streaming or event-based graphs. In contrast, our approach introduces a selective and relevance-aware aggregation mechanism by jointly optimizing node distance and topology. The Topology Reconstructor module adaptively rewires the local structure, filtering out low-utility links and amplifying informative ones across multiple hops.
@@
The concept of heat diffusion has inspired a line of works modeling how signals attenuate over graphs Thanou et al. (2016); Chen et al. (2021); Xu et al. (2020). These models typically employ analytical kernels to encode diffusion patterns into the graph structure. However, they are often static and not integrated with task-driven learning. Our proposed DRTR framework incorporates a diffusion-inspired attenuation mechanism, which adaptively controls information decay based on recomputed distances, enabling fine-grained signal control and improved robustness against over-smoothing or noise amplification in deep GNNs.
@@ 3 Methodology
Unlike standard GNNs that perform shallow 1-hop message passing layer-by-layer, DRTR explicitly models $K$-hop diffusion via direct aggregation across multi-hop neighborhoods. Furthermore, to prevent topological noise propagation, DRTR introduces two key modules:
  • The Distance Recomputator recalibrates semantic distances between nodes in the multi-hop domain and prunes noisy edges.
  • The Topology Reconstructor establishes latent high-quality connections between semantically related but structurally distant nodes.
@@ 3.1 Multi-Hop Diffusion Aggregation
$\alpha_{vu}^{(k)}=\frac{\exp\left(\text{LeakyReLU}(\mathbf{a}^{\top}[\mathbf{W}\mathbf{x}_{v}\,\|\,\mathbf{W}\mathbf{x}_{u}])/\tau_{k}\right)}{\sum_{u^{\prime}\in\mathcal{N}^{(k)}(v)}\exp\left(\text{LeakyReLU}(\mathbf{a}^{\top}[\mathbf{W}\mathbf{x}_{v}\,\|\,\mathbf{W}\mathbf{x}_{u^{\prime}}])/\tau_{k}\right)}.$ (2)
@@
where $\mu_{v}^{(k)}=\frac{1}{d}\sum_{i=1}^{d}h_{v,i}^{(k)}$ and $\sigma_{v}^{(k)}=\sqrt{\frac{1}{d}\sum_{i=1}^{d}(h_{v,i}^{(k)}-\mu_{v}^{(k)})^{2}}$.
@@
where $\{\phi_{k}\}_{k=1}^{K}$ are learnable parameters that control the relative importance of different hop distances.
@@ 3.2 Distance Recomputator (DR)
To mitigate topological noise, we introduce a dynamic Distance Recomputator module that filters weak edges based on learned semantic distance. The pairwise distance between $v$ and $u\in\mathcal{N}^{(k)}(v)$ is defined as:
@@
where $\text{hop\_penalty}(k)=k^{2}$ penalizes longer paths, and $\text{semantic\_divergence}(\mathbf{x}_{v},\mathbf{x}_{u})=1-\cos(\mathbf{x}_{v},\mathbf{x}_{u})$ measures feature dissimilarity.
@@
where the threshold $\alpha_{k}$ is adaptively computed as:
$\alpha_{k}=\mu_{k}+\sigma_{k}\cdot\text{percentile}(p),$ (11)
with $\mu_{k}$ and $\sigma_{k}$ being the mean and standard deviation of distances at hop $k$, and $p$ is a hyperparameter controlling the pruning ratio.
The distance recomputation process is optimized through gradient-based learning:
$\frac{\partial d_{vu}^{(k)}}{\partial\mathbf{x}_{v}}=2(\mathbf{x}_{v}-\mathbf{x}_{u})+\lambda_{k}\frac{\partial\delta^{(k)}_{vu}}{\partial\mathbf{x}_{v}}.$ (12)
@@ 3.3 Topology Reconstructor (TR)
The Topology Reconstructor augments the graph with new edges to improve structural balance. We identify semantically similar but topologically distant nodes based on a multi-criteria similarity function:
$s_{vu}=\omega_{1}\cdot\left\|\mathbf{x}_{v}-\mathbf{x}_{u}\right\|_{2}^{2}+\omega_{2}\cdot\text{structural\_similarity}(v,u)+\omega_{3}\cdot\text{contextual\_alignment}(v,u),$ (13)
where $\omega_{1},\omega_{2},\omega_{3}$ are learnable weights.
@@
where $\sigma(\cdot)$ is the sigmoid function and $\beta$ is the similarity threshold.
@@
The topology reconstruction process is optimized through a joint objective:
$\mathcal{L}_{\text{TR}}=\sum_{(v,u)\in\mathcal{E}_{\text{candidate}}}\left[s_{vu}\cdot\log P(\text{add edge }(v,u))+(1-s_{vu})\cdot\log(1-P(\text{add edge }(v,u)))\right].$ (18)
@@ 3.4 DRTR Training Procedure
Input : Graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, features $\mathbf{X}$, labels $\{y_{v}\}_{v\in\mathcal{V}_{L}}$
Output : Node embeddings $\mathbf{Z}$, prediction scores $\hat{\mathbf{y}}$
1 Initialize parameters $\{\mathbf{W}^{(k)},\gamma_{k}\}_{k=1}^{K}$
2 for each training step $t$ do
 // K-hop Multi-Diffusion Aggregation
3 for each $v\in\mathcal{V}$ do
4    for $k=1$ to $K$ do
5         compute $\mathcal{N}^{(k)}(v)$ via static sampling
6         compute $\mathbf{h}_{v}^{(k)}$ using heat attention (Eq. 2)
7
8    $\mathbf{z}_{v}\leftarrow\sum_{k}\gamma_{k}\mathbf{h}_{v}^{(k)}$
9
 // Distance Recomputator and Topology Update
10   recompute $\{d_{vu}^{(k)}\}$ for all $u\in\mathcal{N}^{(k)}(v)$
11   prune neighbors with $d_{vu}^{(k)}>\alpha$
12   add new edges for $\|\mathbf{x}_{v}-\mathbf{x}_{u}\|_{2}^{2}<\beta$
13   update $\mathcal{G}$ and repeat
14
Algorithm 1 Unified DRTR Training Procedure

### v7

@@ Abstract
We propose DRTR, an efficient framework that integrates distance-aware multi-hop message passing with dynamic topology refinement. Unlike standard GNNs that rely on shallow, fixed-hop aggregation, DRTR leverages both static preprocessing and dynamic resampling to capture deeper structural dependencies. A Distance Recomputator prunes semantically weak edges using adaptive attention, while a Topology Reconstructor establishes latent connections among distant but relevant nodes. This joint mechanism enables more expressive and robust graph representation optimization across evolving graph structures. Extensive experiments demonstrate that DRTR outperforms baseline GNNs in both accuracy and scalability, with at most 20% computational overhead, especially in complex and noisy graph environments.
@@ 1 Introduction
Graph Neural Networks (GNNs) have emerged as a powerful paradigm for semi-supervised node classification and graph representation learning, particularly due to their ability to propagate information over graph structures via message passing kipf2016semi; hamilton2017inductive. Classic GNNs operate by aggregating features from multi-hop neighborhoods using static architectures and fixed aggregation depths. However, such designs may fall short when handling graphs with noisy connections, uneven structural densities, or dynamically evolving topologies Liu2021graphres; wang2021esti, leading to specific failure modes such as over-smoothing in deep layers, noise over-amplification through multi-hop propagation, and hub-dominated aggregation that biases representations toward high-degree nodes.
Recent studies have started exploring adaptive multi-hop aggregation. For example, GAMLP Zhang_2022 applies distinct MLPs for each hop-wise representation, while ImprovingTE Yao2023ImprovingTE enriches hop-specific encodings with contextual substructures. Nevertheless, these methods still depend heavily on fixed-hop sampling strategies, which may hinder generalization and limit the ability to capture complex or noisy graph signals.
In this work, we view graph optimization from a representation-centric perspective: many downstream optimization tasks—such as routing, matching, recommendation, and property prediction—critically rely on the quality of node and edge representations. DRTR optimizes the computational graph by refining multi-hop neighborhoods and reconstructing topology, thereby serving as an efficient optimization layer that improves the solution quality of downstream graph optimization problems. We propose a novel adaptive reconstruction framework for multi-hop GNNs that dynamically adjusts both distance computations and local graph structures to promote more effective and robust message passing. Inspired by coreset selection guo2022deepcore, which selects a small, informative subset of data points to approximate the full dataset, our method adaptively reconstructs the neighborhood for each node by selecting a small, informative subset of neighbors to approximate the full multi-hop neighborhood, enabling the model to better capture latent semantic patterns and mitigate the propagation of noise.
@@
Unlike previous models that statically aggregate neighborhood features, DRTR continuously adapts its perception of graph topology and node relationships throughout training. This dynamic reconstruction mechanism improves the quality of learned representations, stabilizes optimization dynamics, and enables better generalization across a wide range of graphs.
To validate the effectiveness of our framework, we conduct comprehensive experiments across several benchmark datasets, including both static and dynamic graph settings. Our results demonstrate that DRTR consistently outperforms state-of-the-art multi-hop GNN models in terms of accuracy, robustness, and stability. Furthermore, our theoretical analysis sheds light on why adaptive graph refinement enhances graph representation optimization by aligning information flow with the underlying graph semantics.
Our contributions are threefold:
(i) We introduce DRTR, a distance-aware multi-hop framework that explicitly separates diffusion, distance recomputation, and topology reconstruction, enabling robust aggregation under noisy and evolving graphs.
(ii) We design two complementary modules—the Distance Recomputator (DR) and Topology Reconstructor (TR)—which jointly prune noisy edges and add latent long-range links in a task-driven manner, improving representation quality and model robustness in both static and evolving graph scenarios.
(iii) We provide empirical and theoretical analysis showing that DRTR reduces effective neighborhood size and improves generalization: empirically, DRTR consistently improves accuracy and stability across diverse benchmarks and downstream tasks; theoretically, we establish generalization, convergence, and stability guarantees under adaptive topology refinement.
@@ 2 Related Works
Unlike classical combinational graph optimization (e.g., shortest paths, routing, matching), representation-centric graph optimization focuses on optimizing the message-passing pathways and refining graph structure to support downstream optimization tasks. DRTR belongs to this second class, where graph optimization refers to optimizing graph structure, optimizing message passing paths, and improving downstream optimization tasks.
@@
Conventional Graph Neural Networks (GNNs) such as GCN kipf2016semi, GraphSAGE hamilton2017inductive, and GAT vel2018graph adopt fixed-hop message passing schemes and pre-defined neighbor aggregation strategies. However, recent studies recognize that structural connections in graphs often suffer from noise, imbalance, or redundancy liu2023survey. To address this, various works have explored structure learning franceschi2019learning; chen2020iterative; Liu2021graphres, which attempts to refine or reconstruct the graph topology during training. Nonetheless, many such methods rely on end-to-end optimization or learn a static edge mask, which lacks the flexibility to respond to evolving node semantics during multi-hop propagation.
@@
Several studies have attempted to incorporate distance or importance-based filtering into message passing. For instance, Personalized PageRank (PPR)-based methods gasteiger2018pp; chien2021adaptive reweight neighbor contributions based on precomputed transition probabilities. Meanwhile, GAMLP Zhang_2022 and ImprovingTE Yao2023ImprovingTE explicitly leverage multi-hop propagation, with hop-specific encoders and contextual subgraph tokens. However, these approaches still depend on fixed-hop sampling, which may not fully capture varying node proximities or information decay across hops. Unlike prior work, DRTR dynamically recomputes node distances during training using the proposed Distance Recomputator, allowing for more context-aware and adaptive neighborhood construction.
@@
Traditional GNNs typically perform synchronous aggregation, where all neighbor messages are processed simultaneously within each hop wu2020comprehensive. This design overlooks the fact that different neighbors may carry drastically different relevance to the target node. Asynchronous strategies li2017gated; 9878467 attempt to stagger updates or introduce temporal dynamics, but often target streaming or event-based graphs. In contrast, our approach introduces a selective and relevance-aware aggregation mechanism by jointly optimizing node distance and topology. The Topology Reconstructor module adaptively rewires the local structure, filtering out low-utility links and amplifying informative ones across multiple hops.
@@
The concept of heat diffusion has inspired a line of works modeling how signals attenuate over graphs thanou2016learning; Chen_2021_prop; xu2020heatk. These models typically employ analytical kernels to encode diffusion patterns into the graph structure. However, they are often static and not integrated with task-driven learning. Our proposed DRTR framework incorporates a diffusion-inspired attenuation mechanism, which adaptively controls information decay based on recomputed distances, enabling fine-grained signal control and improved robustness against over-smoothing or noise amplification in deep GNNs.
@@ 3 Methodology
Unlike standard GNNs that perform shallow 1-hop message passing layer-by-layer, DRTR explicitly models $K$-hop diffusion via direct aggregation across multi-hop neighborhoods. To prevent topological noise propagation, DRTR introduces two key modules: the Distance Recomputator, which recalibrates semantic distances between nodes in the multi-hop domain and prunes noisy edges, and the Topology Reconstructor, which establishes latent high-quality connections between semantically related but structurally distant nodes.
@@ 3.1 Multi-Hop Diffusion Aggregation
$\alpha_{vu}^{(k)}=\frac{\exp\left(\text{LeakyReLU}(\mathbf{a}^{\top}[\mathbf{W}^{(k)}\mathbf{x}_{v}\,\|\,\mathbf{W}^{(k)}\mathbf{x}_{u}])/\tau_{k}\right)}{\sum_{u^{\prime}\in\mathcal{N}^{(k)}(v)}\exp\left(\text{LeakyReLU}(\mathbf{a}^{\top}[\mathbf{W}^{(k)}\mathbf{x}_{v}\,\|\,\mathbf{W}^{(k)}\mathbf{x}_{u^{\prime}}])/\tau_{k}\right)}.$ (2)
Here, $\mathbf{W}^{(k)}$ is a hop-specific transformation matrix that enables different feature transformations at different hop distances, while $\mathbf{a}$ is a shared attention vector.
@@
$\tilde{\mathbf{h}}_{v}^{(k)}=\text{LayerNorm}\left(\mathbf{h}_{v}^{(k)}\right),$ (4)
which normalizes the activations to have zero mean and unit variance (see Appendix for detailed formulation).
@@
where $\{\phi_{k}\}_{k=1}^{K}$ are learnable parameters that control the relative importance of different hop distances. We use global hop-level weights $\gamma_{k}$ rather than per-node gating (e.g., $\gamma_{k}(v)=\text{softmax}(\bm{\psi}_{k}^{\top}\tilde{\mathbf{h}}_{v}^{(k)})$) to maintain computational efficiency and ensure consistent hop importance across the graph, which is particularly important for distance-aware aggregation where hop semantics should be globally consistent.
@@ 3.2 Distance Recomputator (DR)
This module performs local graph optimization by recalibrating node-to-node distances and removing edges that degrade downstream optimization quality. To mitigate topological noise, we introduce a dynamic Distance Recomputator module that filters weak edges based on learned semantic distance. The pairwise distance between $v$ and $u\in\mathcal{N}^{(k)}(v)$ is defined as:
@@
where $\text{hop\_penalty}(k)=k^{2}$ penalizes longer paths, and $\text{semantic\_divergence}(\mathbf{x}_{v},\mathbf{x}_{u})=1-\cos(\mathbf{x}_{v},\mathbf{x}_{u})$ measures feature dissimilarity. The distance definition combines Euclidean distance $\|\mathbf{x}_{v}-\mathbf{x}_{u}\|_{2}^{2}$ (capturing magnitude differences) with cosine-based semantic divergence (capturing directional differences), where $\beta_{1}$ and $\beta_{2}$ control the trade-off between structural depth penalty and semantic similarity.
@@
where the threshold $\alpha_{k}$ is set as the $p$-th percentile of $\{d_{vu}^{(k)}\}_{u\in\mathcal{N}^{(k)}(v)}$, which adaptively controls the pruning ratio. This percentile-based approach ensures that a consistent fraction of neighbors is retained across different nodes and hops, maintaining stable aggregation patterns.
The distance recomputation process is end-to-end differentiable, enabling gradient-based optimization of the distance metric alongside the main classification objective (see Appendix for gradient details).
@@ 3.3 Topology Reconstructor (TR)
This module performs global topology optimization by adding latent edges that improve connectivity and information flow for downstream optimization tasks. The Topology Reconstructor augments the graph with new edges to improve structural balance. We identify semantically similar but topologically distant nodes based on a multi-criteria similarity function. To maintain consistency (where larger values indicate greater similarity), we define:
$\text{sim}_{vu}=\omega_{1}\cdot\text{contextual\_alignment}(v,u)+\omega_{2}\cdot\text{structural\_similarity}(v,u)-\omega_{3}\cdot\left\|\mathbf{x}_{v}-\mathbf{x}_{u}\right\|_{2}^{2},$ (11)
where $\omega_{1},\omega_{2},\omega_{3}\geq 0$ are learnable weights.
@@
where $\sigma(\cdot)$ is the sigmoid function, $\beta$ is the similarity threshold, and $\tau$ is a temperature parameter controlling the sharpness of the probability distribution.
@@
To ensure computational efficiency, we restrict candidate edge selection to approximate $k$-nearest neighbors in the embedding space, limiting the candidate set $\mathcal{E}_{\text{candidate}}$ to at most $R$ edges per node. This reduces the complexity from $\mathcal{O}(|\mathcal{V}|^{2})$ to $\mathcal{O}(|\mathcal{V}|\cdot k\cdot d)$ where $k$ is the number of approximate nearest neighbors and $d$ is the feature dimension.
The topology reconstruction process is optimized through a margin-based objective that encourages high-similarity pairs to be connected:
$\mathcal{L}_{\text{TR}}=\sum_{(v,u)\in\mathcal{E}_{\text{candidate}}}\max(0,\beta-\text{sim}_{vu})\cdot P(\text{add edge }(v,u)),$ (16)
which penalizes high-similarity pairs that are not connected, encouraging the model to add edges between semantically similar nodes.
@@ 3.4 DRTR Training Procedure
[…]
Output: Node embeddings $\mathbf{Z}\in\mathbb{R}^{N\times d^{\prime}}$
Initialize $Q_{h}^{(k)}(v)\leftarrow 0$, $N_{h}^{(k)}(v)\leftarrow 0$ for all $(v,k,h)\in\mathcal{V}\times[K]\times[K]$;
Initialize $\mathbf{W}^{(k)}\in\mathbb{R}^{d\times d^{\prime}}$, $\phi_{k}$ for all $k\in[K]$, $\gamma_{k}\leftarrow\text{softmax}(\phi_{k})$;
for training iteration $t=1,\ldots,T$ do
 for each node $v\in\mathcal{V}$ do
    for $k=1,\ldots,K$ do
       $\mathcal{N}^{(k)}(v)\leftarrow\text{StaticSampling}(\mathcal{G},v,k)$;
         Compute $\alpha_{vu}^{(k)}$ via Eq. (2);
       $\mathbf{h}_{v}^{(k)}\leftarrow\sum_{u\in\mathcal{N}^{(k)}(v)}\alpha_{vu}^{(k)}\cdot\mathbf{W}^{(k)}\mathbf{x}_{u}$;
       $\tilde{\mathbf{h}}_{v}^{(k)}\leftarrow\text{LayerNorm}(\mathbf{h}_{v}^{(k)})$;
      end for
    $\mathbf{z}_{v}\leftarrow\sum_{k=1}^{K}\gamma_{k}\cdot\tilde{\mathbf{h}}_{v}^{(k)}$;
   end for
 for each node $v\in\mathcal{V}$ do
    for $k=1,\ldots,K$ do
       for each $u\in\mathcal{N}^{(k)}(v)$ do
          $n\leftarrow N_{k}(v,u)\leftarrow N_{k}(v,u)+1$;
          $d_{vu}^{(k)}\leftarrow\|\mathbf{x}_{v}-\mathbf{x}_{u}\|_{2}^{2}+\lambda_{k}\cdot\delta_{vu}^{(k)}$;
         end for
       $\alpha_{k}\leftarrow p\text{-th percentile of }\{d_{vu}^{(k)}\}_{u\in\mathcal{N}^{(k)}(v)}$;
       $\mathcal{N}^{(k)}(v)\leftarrow\{u\in\mathcal{N}^{(k)}(v)\mid d_{vu}^{(k)}\leq\alpha_{k}\}$;
      end for
   end for
 $\mathcal{E}_{\text{candidate}}\leftarrow\text{ApproximateKNN}(\mathbf{X},k)$;
 for each $(v,u)\in\mathcal{E}_{\text{candidate}}\setminus\mathcal{E}$ do
    $\text{sim}_{vu}\leftarrow\omega_{1}\cdot\text{contextual\_alignment}(v,u)+\omega_{2}\cdot\text{structural\_similarity}(v,u)-\omega_{3}\cdot\|\mathbf{x}_{v}-\mathbf{x}_{u}\|_{2}^{2}$;
    if $\sigma((\text{sim}_{vu}-\beta)/\tau)>\theta$ then
       $\mathcal{E}\leftarrow\mathcal{E}\cup\{(v,u)\}$;
      end if
   end for
  Compute loss $\mathcal{L}$ and update parameters via gradient descent;
end for
Algorithm 1 DRTR Training Algorithm